# Haoran Guan

### v1

@@ 3.3.1 Orthogonality
For $\|E_{2}\|_{2}$, we put (37), (42), (46), (50), (53) and (56) into (35). With (21), (23) and (24), we can have
$\|E_{2}\|_{2}$$\leq$ $\|UR^{-1}\|_{2}^{2}\|E_{HQ}\|_{2}^{2}+2\|UR^{-1}\|_{2}\|GUR^{-1}\|_{2}\|E_{HQ}\|_{2}$ (65)
@@
$\leq$ $(\frac{1.29}{\sigma_{min}(L)}\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2})^{2}+2\cdot\frac{1.29}{\sigma_{min}(L)}\cdot 1.03\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2}$
$+$ $2\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$
@@
For $\|E_{3}\|_{2}$, we put (39) and (42) into (36). With (24), we can get
$\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}$ (66)
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$
@@
Therefore, we put (64)-(66) into (30) and we can get
$\|Q^{\top}Q-I\|_{2}\leq 0.74.$ (67)
With (67), we can get
$\|Q\|_{2}$$\leq$ $\sqrt{1+0.74}$ (68)
@@
Based on (68) and (69), we can get
$\kappa_{2}(Q)$$\leq$ $\frac{\|Q\|_{2}}{\sigma_{min}(Q)}$ (70)
@@
According to the theoretical result in error , for LHC2 with a step of CholeskyQR after LHC, based on (70), we can get
@@
(25) is proved.
3.3.2 Residual
In the following, we consider the residual. We write CholeskyQR after LHC with error matrices below.
$B=Q^{\top}Q+E_{A},$ (71)
$R_{1}^{\top}R_{1}=B+E_{B},$ (72)
$Q_{1}R_{1}=Q+E_{C},$ (73)
$R_{2}=R_{1}R+E_{D}.$ (74)
Based on (73) and (74), we can have
$Q_{1}R_{2}$ $=$ $(Q+E_{C})R_{1}^{-1}(R_{1}R+E_{D})$ (75)
@@ 3.3.2 Residual
With (75), we can get
$\|Q_{1}R_{2}-X\|_{F}\leq\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}.$ (76)
With Lemma 2, we can have
$\|E_{C}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2},$ (77)
$\|E_{D}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{1}\|_{2}\|R\|_{2},$ (78)
$\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}.$ (79)
We put (77)-(79) into (76) and we can get
$\|QR-X\|_{F}$$\leq$ $\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}$ (80)
$\leq$ $1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}$
$\leq$ $1.1n^{2}{\bf u}(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}).$
With (21), (22), (25) and the theoretical result in error , we can get
$\|Q_{1}\|_{2}$$\leq$ $\sqrt{1+34.85\cdot(\frac{1}{64}+\frac{1}{64})}$ (81)
@@
With Lemma 1, (20), (79) and properties of singular values, we can have
$\|X\|_{2}$ $=$ $\|QR-E_{QR}\|_{2}$ (83)
$\geq$ $\sigma_{min}(QR-E_{QR})$
$\geq$ $\sigma_{min}(Q)\|R\|_{2}-1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}.$
With (22), (68), (69) and (83), we can have
$\|R\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\sigma_{min}(Q)-1.1n^{2}{\bf u}\|Q\|_{2}}$ (84)
@@
Therefore, we put (68), (81), (82) and (84) into (80) and we can have
$\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2})$
@@
$\leq$ $7.94n^{2}{\bf u}\|X\|_{2}.$
(26) is proved. In all, Theorem 3.1 is proved.
4 Randomized LU-Householder CholeskyQR
@@
4.1 Theoretical results of SSLHC3 and general settings
In this part, we analyze SSLHC3. We write SSLHC first with error matrices below:
$PX$ $=$ $LU+E_{LU},$ (85)
$L_{ss}$ $=$ $\Omega_{2}\Omega_{1}L+E_{Lss},$ (86)
$L_{ss}$ $=$ $W_{ss}G_{ss}+E_{HQss},$ (87)
$R_{ss}$ $=$ $G_{ss}U-E_{Rss},$ (88)
$X$ $=$ $Q_{ss}R_{ss}-E_{QRss},$ (89)
$B_{1ss}$ $=$ $Q_{ss}^{\top}Q_{ss}+E_{1ss},$ (90)
$R_{1ss}^{\top}R_{1ss}$ $=$ $B_{1ss}+E_{2ss},$ (91)
$Q_{1ss}R_{1ss}$ $=$ $Q_{ss}+E_{3ss},$ (92)
$R_{2ss}$ $=$ $R_{1ss}R_{ss}+E_{4ss},$ (93)
$B_{2ss}$ $=$ $Q_{1ss}^{\top}{Q_{2ss}}+E_{5ss},$ (94)
$R_{3ss}^{\top}R_{3ss}$ $=$ $B_{2ss}+E_{6ss},$ (95)
$Q_{2ss}R_{3ss}$ $=$ $Q_{1ss}+E_{7ss},$ (96)
$R_{4ss}$ $=$ $R_{3ss}R_{2ss}+E_{8ss}.$ (97)
@@ 4.1 Theoretical results of SSLHC3 and general settings
To analyze SSLHC with the multi-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding. We provide the following assumption.
Assumption 1 (The assumption for the multi-sketching)
@@
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, we can have
$\frac{\sqrt{1-\epsilon_{s}}}{\sqrt{1+\epsilon_{b}}}>\max(12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1,0.14)$ (98)
with probability at least $1-p_{f}$. Here, we define $p_{f}=p_{1}+p_{2}-p_{1}p_{2}$, $\epsilon_{s}=\epsilon_{1}+\epsilon_{2}-\epsilon_{1}\epsilon_{2}$ and $\epsilon_{b}=\epsilon_{1}+\epsilon_{2}+\epsilon_{1}\epsilon_{2}$.
@@
With Assumption 1, we provide the primary theorem of rounding error analysis of SSLHC3.
Theorem 4.1 (Rounding error analysis of SSLHC3)
@@
Suppose Assumption 1 is satisfied, if $X\in\mathbb{R}^{m\times n}$, when $[Q_{2ss},R_{4ss}]=SSLHC3(X)$, we can have
$\|Q_{2ss}^{\top}Q_{2ss}-I\|_{F}$$\leq$ $6(mn{\bf u}+n(n+1){\bf u}),$ (99)
$\|Q_{2ss}R_{4ss}-X\|_{F}$$\leq$ $R$ (100)
@@
with probability at least $1-p_{f}$. Here, $R=[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\|X\|_{2}$, $h_{1}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{2}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}}$.
For SSLHC3, the following assumptions hold:
$s_{1}\sqrt{n}{\bf u}\leq\frac{1}{64},$ (101)
$s_{2}n^{2}{\bf u}\leq\frac{1}{64},$ (102)
$24.2\cdot\frac{\sqrt{1+\epsilon_{b}}}{\sqrt{1-\epsilon_{s}}}\cdot\sqrt{n}{\bf u}\cdot(s_{1}+s_{2}n\sqrt{n})\kappa_{2}(L)$$\leq$ $1,$ (103)
$22.6n^{2}\sqrt{n}{\bf u}\cdot\frac{\sqrt{1+\epsilon_{b}}}{\sqrt{1-\epsilon_{s}}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 . (104)
4.2 Theoretical results of SLHC2 and general settings
@@
To analyze SLHC with the single-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding. We provide the following assumption.
Assumption 2 (The assumption for the single-sketching)
If $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we can have
$\frac{\sqrt{1-\epsilon}}{\sqrt{1+\epsilon}}>12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1$ (105)
with probability at least $1-p$.
With Assumption 2, we provide the primary theorem of rounding error analysis of SLHC2.
Theorem 4.2 (Rounding error analysis of SLHC2)
Suppose Assumption 2 is satisfied, if $X\in\mathbb{R}^{m\times n}$, when $[Q_{1ss},R_{2ss}]=SLHC2(X)$, we can have
$\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $h_{3},$ (106)
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $T$ (107)
with probability at least $1-p_{f}$. Here, $T=(2.93\sqrt{1+h_{3}}+1.41)\cdot\frac{h_{4}}{\sqrt{1-\epsilon}}\cdot n^{2}{\bf u}\|X\|_{2}$, $h_{3}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon}{1+\epsilon}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{4}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon}}-\frac{0.11}{\sqrt{1-\epsilon}}}$.
@@
$sn^{2}{\bf u}\leq\frac{1}{64},$ (108)
$24.2\cdot\frac{\sqrt{1+\epsilon}}{\sqrt{1-\epsilon}}\cdot\sqrt{n}{\bf u}\cdot(m+sn\sqrt{n})\kappa_{2}(L)$$\leq$ $1,$ (109)
$22.6n^{2}\sqrt{n}{\bf u}\cdot\frac{\sqrt{1+\epsilon}}{\sqrt{1-\epsilon}}\cdot(1+1.08sn\sqrt{n}{\bf u})\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 . (110)
4.3 Preparations for proving Theorem 4.1
@@
Lemma 17 (Estimation of $\|E_{Lss}\|_{2}$)
For $\|E_{Lss}\|_{2}$, we can have
$\|E_{Lss}\|_{2}\leq 1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$ (111)
@@
Proof
To estimate $\|E_{Lss}\|_{2}$, we divide (86) into two steps.
$L_{t}=\Omega_{1}L+E_{L_{1}},$ (112)
$L_{ss}=\Omega_{2}L_{t}+E_{L_{2}}.$ (113)
When we make comparison between (112), (113) and (86), we can have
$E_{Lss}=\Omega_{2}E_{L_{1}}+E_{L_{2}}.$ (114)
Based on Lemma 2 and Lemma 5, we can have
$\|E_{L_{1}}\|_{2}$$\leq$ $\|E_{L_{1}}\|_{F}$ (115)
$\leq$ $\gamma_{m}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{1}$. Therefore, with Lemma 2 and (115), we can have
$\|E_{L_{2}}\|_{2}$$\leq$ $\|E_{L_{2}}\|_{F}$ (116)
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot\|\Omega_{1}L+E_{L_{1}}\|_{F}$
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot(\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}+\gamma_{m}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2})$
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{(1+\epsilon_{1})(1+\epsilon_{2})}\cdot(1+\gamma_{m})\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. With (115), we can have
$\|\Omega_{2}E_{L_{1}}\|_{2}$$\leq$ $\|\Omega_{2}E_{L_{1}}\|_{F}$ (117)
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot\|E_{L_{1}}\|_{F}$
$\leq$ $\gamma_{s_{1}}\cdot\gamma_{m}\cdot\sqrt{(1+\epsilon_{1})(1+\epsilon_{2})}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. Under the definition in Lemma 7, we put (116) and (117) into (114) and we can have

### v2

@@ 3.3.1. Orthogonality
For $\|E_{2}\|_{2}$, we put (3.21), (3.26), (3.30), (3.34), (3.37) and (3.40) into (3.19) and with (3.5), (3.7) and (3.8), we can have
(3.49) $\|E_{2}\|_{2}$$\leq$ $\|UR^{-1}\|_{2}^{2}\|E_{HQ}\|_{2}^{2}+2\|UR^{-1}\|_{2}\|GUR^{-1}\|_{2}\|E_{HQ}\|_{2}$
@@
$\leq$ $(\frac{1.29}{\sigma_{min}(L)}\cdot 1.1mn\sqrt{n}{\bf u}\cdot\|L\|_{2})^{2}+2\cdot\frac{1.29}{\sigma_{min}(L)}\cdot 1.03\cdot 1.1mn\sqrt{n}{\bf u}\cdot\|L\|_{2}$
$+$ $2\cdot 1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2})^{2}$
@@
For $\|E_{3}\|_{2}$, we put (3.23) and (3.26) into (3.20) and with (3.8), we can get
(3.50) $\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}$
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2})^{2}$
@@
Therefore, we put (3.48)-(3.50) into (3.14) and we can get
(3.51) $\|Q^{\top}Q-I\|_{2}\leq 0.74.$
With (3.51), it is obvious to see that
(3.52) $\|Q\|_{2}$$\leq$ $\sqrt{1+0.74}$
@@
Based on (3.52) and (3.53), we can get
(3.54) $\kappa_{2}(Q)$$\leq$ $\frac{\|Q\|_{2}}{\sigma_{min}(Q)}$
@@
According to the theoretical result in [39], for LHC2 with a step of CholeskyQR after LHC, based on (3.54), we can get
@@
(3.9) is proved.
3.3.2. Residual
In the following, we consider the residual. We write CholeskyQR after LHC with error matrices below:
(3.55) $B=Q^{\top}Q+E_{A},$
(3.56) $R_{1}^{\top}R_{1}=B+E_{B},$
(3.57) $Q_{1}R_{1}=Q+E_{C},$
(3.58) $R_{2}=R_{1}R+E_{D}.$
Based on (3.57) and (3.58), we can have
(3.59) $Q_{1}R_{2}$ $=$ $(Q+E_{C})R_{1}^{-1}(R_{1}R+E_{D})$
@@ 3.3.2. Residual
With (3.59), we can get
(3.60) $\|Q_{1}R_{2}-X\|_{F}\leq\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}.$
Based on Lemma 2.2 and similar to the previous results, we can have
(3.61) $\|E_{C}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\cdot\|Q_{1}\|_{2}\|R_{1}\|_{2},$
(3.62) $\|E_{D}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\cdot\|R_{1}\|_{2}\|R\|_{2},$
(3.63) $\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}\cdot\|Q\|_{2}\|R\|_{2}.$
We put (3.61)-(3.63) into (3.60) and we can get
(3.64) $\|QR-X\|_{F}$$\leq$ $\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}$
$\leq$ $1.1n^{2}{\bf u}\cdot\|Q\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\cdot\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\cdot\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}$
$\leq$ $1.1n^{2}{\bf u}\cdot(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}).$
With (3.5), (3.6), (3.9) and the theoretical result in [39], we can get
(3.65) $\|Q_{1}\|_{2}$$\leq$ $\sqrt{1+34.85\cdot(\frac{1}{64}+\frac{1}{64})}$
@@
With Lemma 2.1, (3.4), (3.63) and properties of singular values, we can have
(3.67) $\|X\|_{2}$ $=$ $\|QR-E_{QR}\|_{2}$
$\geq$ $\sigma_{min}(Q)\|R\|_{2}-1.1n^{2}{\bf u}\cdot\|Q\|_{2}\|R\|_{2}.$
With (3.6), (3.52), (3.53) and (3.67), we can have
(3.68) $\|R\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\sigma_{min}(Q)-1.1n^{2}{\bf u}\cdot\|Q\|_{2}}$
@@
Therefore, we put (3.52), (3.65), (3.66) and (3.68) into (3.64) and we can have
$\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}\cdot(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2})$
@@
$\leq$ $7.94n^{2}{\bf u}\cdot\|X\|_{2}.$
(3.10) is proved. In all, Theorem 3.1 is proved. ∎
4. Randomized LU-Householder CholeskyQR
@@
4.1. Theoretical results of SSLHC3 and general settings
In this part, we analyze SSLHC3. We write SSLHC3 first with error matrices below:
(4.1) $PX$ $=$ $LU+E_{LU},$
(4.2) $L_{ss}$ $=$ $\Omega_{2}\Omega_{1}L+E_{Lss},$
(4.3) $L_{ss}$ $=$ $W_{ss}G_{ss}+E_{HQss},$
(4.4) $R_{ss}$ $=$ $G_{ss}U-E_{Rss},$
(4.5) $X$ $=$ $Q_{ss}R_{ss}-E_{QRss},$
$B_{1ss}$ $=$ $Q_{ss}^{\top}Q_{ss}+E_{1ss},$
$R_{1ss}^{\top}R_{1ss}$ $=$ $B_{1ss}+E_{2ss},$
(4.6) $Q_{1ss}R_{1ss}$ $=$ $Q_{ss}+E_{3ss},$
(4.7) $R_{2ss}$ $=$ $R_{1ss}R_{ss}+E_{4ss},$
$B_{2ss}$ $=$ $Q_{1ss}^{\top}{Q_{2ss}}+E_{5ss},$
$R_{3ss}^{\top}R_{3ss}$ $=$ $B_{2ss}+E_{6ss},$
(4.8) $Q_{2ss}R_{3ss}$ $=$ $Q_{1ss}+E_{7ss},$
(4.9) $R_{4ss}$ $=$ $R_{3ss}R_{2ss}+E_{8ss}.$
@@ 4.1. Theoretical results of SSLHC3 and general settings
To analyze SSLHC with the multi-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in Definition 2.8. We provide the following assumption.
Assumption 1 (The assumption for the multi-sketching).
@@
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, we let
(4.10) $\frac{\sqrt{1-\epsilon_{s}}}{\sqrt{1+\epsilon_{b}}}>\max(12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1,0.14).$
Here, we define $p_{f}=p_{1}+p_{2}-p_{1}p_{2}$, $\epsilon_{s}=\epsilon_{1}+\epsilon_{2}-\epsilon_{1}\epsilon_{2}$ and $\epsilon_{b}=\epsilon_{1}+\epsilon_{2}+\epsilon_{1}\epsilon_{2}$.
For SSLHC3, the following settings hold:
(4.11) $s_{1}\sqrt{s_{2}}{\bf u}\leq\frac{1}{64},$
(4.12) $s_{2}n\sqrt{n}{\bf u}\leq\frac{1}{64},$
(4.13) $\frac{1}{\sqrt{1-\epsilon_{s}}}\cdot(20.4t+22s_{2}n\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}})\cdot\kappa_{2}(L)$$\leq$ $1,$
(4.14) $22.6bn^{2}{\bf u}\cdot\frac{\sqrt{1+\epsilon_{b}}+t}{\sqrt{1-\epsilon_{s}}}\cdot\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 .
Here, $t=(1.1s_{1}{\bf u}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{s_{2}n}\|\Omega_{2}\|_{2}+1.2m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2})$ and $b=\max(\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2},1)$.
Regarding $\|\Omega_{1}\|_{2}$ and $\|\Omega_{2}\|_{2}$, when $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a sketch matrix of CountSketch, there is a deterministic upper bound for $\|\Omega_{1}\|_{2}$ [2, 36]:
$\|\Omega_{1}\|_{2}\leq\sqrt{m}.$
When $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a sketch matrix of Gaussian sketch, a probabilistic bound for $\|\Omega_{2}\|_{2}$ holds
$\|\Omega_{2}\|_{2}\leq 1+C(\sqrt{\frac{s_{1}}{s_{2}}}+\frac{3}{\sqrt{s_{2}}})$
[…]
@@
With Assumption 1 and the general settings above, we provide the primary theorem of rounding error analysis of SSLHC3.
Theorem 4.1 (Rounding error analysis of SSLHC3).
@@
Suppose Assumption 1 and (4.11)-(4.14) are satisfied, for $X\in\mathbb{R}^{m\times n}$ and $[Q_{2ss},R_{4ss}]=\mbox{SSLHC3}(X)$, we have
(4.15) $\|Q_{2ss}^{\top}Q_{2ss}-I\|_{F}$$\leq$ $6(mn{\bf u}+n(n+1){\bf u}),$
(4.16) $\|Q_{2ss}R_{4ss}-X\|_{F}$$\leq$ $R$
@@
with probability at least $(1-p_{f})^{2}$. Here, $R=[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\cdot\|X\|_{2}$, $h_{1}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{2}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}}$.
4.2. Theoretical results of SLHC2 and general settings
@@
To analyze SLHC with the single-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding. With Definition 2.8, we provide the following assumption.
Assumption 2 (The assumption for the single-sketching).
If $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we let
(4.17) $\frac{\sqrt{1-\epsilon}}{\sqrt{1+\epsilon}}>12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1.$
@@
(4.18) $m\sqrt{s}{\bf u}\leq\frac{1}{64},$
(4.19) $sn\sqrt{n}{\bf u}\leq\frac{1}{64},$
(4.20) $\frac{1}{\sqrt{1-\epsilon}}\cdot(20.4j+22s_{2}n\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon})\cdot\kappa_{2}(L)$$\leq$ $1,$
(4.21) $22.6dn^{2}{\bf u}\cdot\frac{\sqrt{1+\epsilon}+j}{\sqrt{1-\epsilon}}\cdot\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 .
Here, $j=(1.1m{\bf u}\cdot\sqrt{s_{2}n}\|\Omega\|_{2}+1.2m{\bf u}\cdot\sqrt{mn}\|\Omega\|_{2})$ and $d=\max(\|\Omega_{2}\|_{2},1)$.
With Assumption 2 and the general settings above, we provide the primary theorem of rounding error analysis of SLHC2.
Theorem 4.2 (Rounding error analysis of SLHC2).
Suppose Assumption 2 and (4.18)-(4.21) are satisfied, for $X\in\mathbb{R}^{m\times n}$ and $[Q_{1ss},R_{2ss}]=\mbox{SLHC2}(X)$, we have
(4.22) $\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $h_{3},$
(4.23) $\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $T$
with probability at least $(1-p)^{2}$. Here, $T=(2.93\sqrt{1+h_{3}}+1.41)\cdot\frac{h_{4}}{\sqrt{1-\epsilon}}\cdot n^{2}{\bf u}\cdot\|X\|_{2}$, $h_{3}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon}{1+\epsilon}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{4}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon}}-\frac{0.11}{\sqrt{1-\epsilon}}}$.
4.3. Preparations for proving Theorem 4.1
@@
Lemma 4.3 (Estimation of $\|E_{Lss}\|_{2}$).
For $\|E_{Lss}\|_{2}$, we have
(4.24) $\|E_{Lss}\|_{2}\leq t\|L\|_{2}$
@@
Proof.
To estimate $\|E_{Lss}\|_{2}$, we rewrite (4.2) with two steps:
(4.25) $L_{t}=\Omega_{1}L+E_{L_{1}},$
(4.26) $L_{ss}=\Omega_{2}L_{t}+E_{L_{2}}.$
With (4.2), (4.25) and (4.26), we can have
(4.27) $E_{Lss}=\Omega_{2}E_{L_{1}}+E_{L_{2}}.$
Based on Lemma 2.2, we can have
(4.28) $\|E_{L_{1}}\|_{2}$$\leq$ $\|E_{L_{1}}\|_{F}$
$\leq$ $1.1m{\bf u}\cdot\|\Omega_{1}\|_{F}\|L\|_{F}$
$\leq$ $1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|L\|_{2}$
with probability at least $1-p_{1}$. Therefore, with Lemma 2.2 and (4.28), we can have
(4.29) $\|E_{L_{2}}\|_{2}$$\leq$ $\|E_{L_{2}}\|_{F}$
$\leq$ $1.1s_{1}{\bf u}\cdot\|\Omega_{2}\|_{F}\|L_{t}\|_{F}$
$\leq$ $1.1s_{1}{\bf u}\cdot\sqrt{s_{2}}\|\Omega_{2}\|_{2}\cdot(\|\Omega_{1}L\|_{F}+\|E_{L_{1}}\|_{F})$
$\leq$ $1.1s_{1}{\bf u}\cdot\sqrt{s_{2}}\|\Omega_{2}\|_{2}\cdot(\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}+1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|L\|_{2})$
with probability at least $1-p_{f}$. With (4.28), we can have
(4.30) $\|\Omega_{2}E_{L_{1}}\|_{2}$$\leq$ $\|\Omega_{2}E_{L_{1}}\|_{F}$
$\leq$ $\|\Omega_{2}\|_{2}\|E_{L_{1}}\|_{F}$
$\leq$ $1.1m{\bf u}\cdot\sqrt{s_{1}n}\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2}\|L\|_{2}$
with probability at least $1-p_{f}$. We put (4.29) and (4.30) into (4.27) and based on (4.11), we can have